OpenStudy (anonymous):

1/3 to the second power over the denominator

jimthompson5910 (jim_thompson5910):

\[\large \frac{1}{3^2}=\frac{1}{9}\]

OpenStudy (anonymous):

sorry thanks .

jimthompson5910 (jim_thompson5910):

Note: \[\large 3^2=3*3=9\]

jimthompson5910 (jim_thompson5910):

sorry about what?

OpenStudy (anonymous):

using your time on a dumb question

jimthompson5910 (jim_thompson5910):

no it's quite alright, it's not a dumb question